# Open-vocabulary Multimodal Emotion Recognition: Dataset, Metric, and Benchmark

### v1

@@ 4.2 Metric Calculation
5 Results and Discussion
In this section, we report the experimental results and provide an in-depth analysis. We default to using GPT-based grouping and employ GPT-3.5 (“gpt-3.5-turbo-16k-0613”) as LLM. We generally report evaluation results in both languages, but if no specific language is mentioned, we default to reporting results for the English branch. To mitigate the impact of randomness, we conduct each experiment twice and report the average scores and standard deviations.
Main Results on CLUE-M/A/T/V.
Table 2 shows the performance of CLUE-M/A/T/V. We observe that CLUE-Multi performs the best. This is because emotions are conveyed through various modalities, and integrating multimodal clues can better understand human emotions. In contrast, CLUE-Text performs the worst. This is because, in MER, text is emotionally ambiguous; without integrating other modalities, we cannot accurately understand the emotions conveyed through text.
Main Results on CLUE-MLLM.
Table 2 presents the results of CLUE-MLLM. Additionally, we introduce a heuristic baseline called Random, where we randomly select a label from basic emotions. This baseline reflects the lower bound of performance. We observe that MLLM generally outperforms Random, indicating that MLLM can partially address OV-MER. However, the performance of MLLM remains unsatisfactory, highlighting the limitations of existing MLLMs and the challenges of OV-MER. Furthermore, models that perform well in Chinese often perform well in English. These results suggest that the impact of language differences on rankings is limited.
Table 3: Performance of combinations of different MLLMs.
Model English Chinese
$\mbox{F}_{\mbox{s}}\uparrow$ $\mbox{Precision}_{\mbox{s}}\uparrow$ $\mbox{Recall}_{\mbox{s}}\uparrow$ $\mbox{F}_{\mbox{s}}\uparrow$ $\mbox{Precision}_{\mbox{s}}\uparrow$ $\mbox{Recall}_{\mbox{s}}\uparrow$
Audio + Text
SECap 45.72±0.09 54.52±0.15 39.37±0.05 45.57±0.13 55.55±0.23 38.64±0.08
SALMONN 47.96±0.04 50.20±0.04 45.92±0.04 48.24±0.03 52.24±0.00 44.82±0.05
Video + Text
Video-ChatGPT 50.52±0.06 54.03±0.04 47.44±0.07 54.73±0.00 61.15±0.10 49.52±0.06
mPLUG-Owl 52.73±0.13 54.54±0.13 51.04±0.13 50.95±0.06 56.40±0.11 46.47±0.18
Chat-UniVi 53.08±0.01 53.68±0.00 52.50±0.02 53.86±0.02 58.54±0.01 49.86±0.03
Audio + Video + Text
SECap + mPLUG-Owl 56.69±0.03 50.05±0.23 65.38±0.33 54.99±0.23 51.65±0.27 58.79±0.16
SECap + Video-ChatGPT 56.90±0.08 52.03±0.04 62.79±0.14 56.49±0.02 56.50±0.01 56.48±0.05
SECap + Chat-UniVi 57.34±0.16 48.85±0.29 69.41±0.13 56.19±0.13 52.38±0.07 60.59±0.22
SALMONN + Video-ChatGPT 58.19±0.23 53.16±0.17 64.26±0.31 55.05±0.16 53.44±0.14 56.76±0.19
SALMONN + Chat-UniVi 58.43±0.06 51.62±0.00 67.31±0.15 56.93±0.06 51.65±0.06 63.42±0.06
SALMONN + mPLUG-Owl 58.70±0.04 51.77±0.01 67.76±0.11 55.62±0.21 51.74±0.19 60.14±0.23
(a) Clue length (H)
(b) Label count (H)
(c) Word cloud (H)
(d) Clue length (H+L)
(e) Label count (H+L)
(f) Word cloud (H+L)
Figure 5: Human-only annotation (H) vs. Human-LLM (H+L) collaboration.
@@ 5 Results and Discussion
Effectiveness of Multimodal Fusion.
In Table 3, we select the best-performing ALLMs and VLLMs and explore whether their combinations can lead to better performance. We observe that these combinations bring performance improvements, with some even surpassing GPT-4V. Therefore, we can more effectively address OV-MER by leveraging the strengths of different MLLMs.
Figure 7: Different ways for generating CLUE-MLLM.
@@
This paper proposes a human-LLM collaborative strategy. In this section, we discuss the necessity of this strategy and test the performance of human-only annotation. In Figure 5, we compare different strategies from three aspects: the length distribution of generated clues, the distribution of label counts, and the word cloud. We observe that the descriptions generated through human-LLM collaboration are longer, indicating that human-only annotations often focus on primary clues while neglecting minor ones. With the pre-annotation and semantic reasoning capabilities of LLMs, we can generate richer and more comprehensive descriptions. Additionally, human-LLM collaboration can produce a broader range of emotions and provide more diverse labels for each sample. These results validate the effectiveness of our strategy.
@@
This section reveals the impact of different CLUE-MLLM generation strategies. Figure 7 introduces three methods: 1) S0 does not use text and inputs the video into MLLM; 2) S1 inputs both text and video into MLLM simultaneously; 3) S2 first uses MLLM to extract descriptions and then combines with text using another LLM, consistent with the strategy shown in Figure 4. In Figure 6, we present the performance of different strategies. Generally, S1 and S2 outperform S0, indicating the importance of text content in OV-MER. Moreover, S2 typically performs better than S1. The reason is that inputting both video and text into the MLLM simultaneously increases the task difficulty, and current MLLMs may struggle to handle complex prompts. S2 divides this process into two steps, reducing task complexity and achieving better performance. Therefore, we adopt S2 as the default strategy.
GPT-based vs. Matching-based Metrics.
In Table 2, CLUE-Multi performs the best. This leads to a hypothesis: Do sentences that are more “similar” to CLUE-Multi yield better emotion recognition performance? The most common way to measure “similarity” is through matching-based metrics, with BLEU11{1}1, BLEU44{4}4, METEOR, and ROUGE1 being the most widely used. Therefore, we use CLUE-MLLM as input and calculate both GPT-based and matching-based metrics (see Figure 8(a)), followed by calculating their PCC scores (see Figure 8(b)). From the experimental results, we have some interesting observations. First, the same metric across different languages typically exhibits high correlations. For example, “F(C)” and “F(E)” have a high PCC score. However, the correlation between GPT-based and matching-based metrics is not strong. For example, the highest PCC score between “F(E)” and matching-based metrics is 0.77. Therefore, a high score in matching-based metrics does not necessarily indicate good performance in OV-MER.
@@
This paper proposes two grouping strategies: GPT-based grouping and EW-based grouping. In this section, we explore the relationship between them and analyze whether the EW-based method can replace the GPT-based method. We calculate the scores for different types of EW-based grouping strategies, with the average score denoted as M-avg. Table 4 reports $\mbox{F}_{\mbox{s}}$, as this metric is used for the final ranking, and we also compute the PCC scores between different metrics. We observe that the PCC score between M-avg and GPT is relatively high, indicating that the EW-based method can replace GPT.
@@
Model GPT M1 M2 M3-W1 M3-W2
L1 L2 L1 L2
PCC score 0.887 0.857 0.911 0.940 0.913 0.942
Model M3-W3 M3-W4 M3-W5 M-avg
L1 L2 L1 L2 L1 L2
PCC score 0.904 0.927 0.899 0.922 0.885 0.894 0.942
6 Discussion
Limitations.
The main contribution of this paper is the definition of a new task and the conduct of foundational research, including the development of datasets, metrics, and solutions. In the future, we plan to design more effective frameworks to solve OV-MER. Meanwhile, we have evaluated some representative MLLMs, but not all models are covered. In the future, we will broaden the scope of our evaluations. In addition, although the OV-MERD dataset is currently small in scale, we intend to expand its size in the future. In fact, preliminary experiments on the OV-MERD dataset provide guidance for our further improvements.
Label Importance.
In OV-MER, we do not assign different levels of importance to each label. We believe that every emotion holds equal significance, and neglecting anyone can impact the performance of downstream tasks. For example, if a human-computer interaction system only captures basic emotions while overlooking nuanced ones, it may fail to generate appropriate responses.
@@ 7 Conclusion
Reproducibility Statement
This paper provides the source code and intermediate results in the supplementary materials. Since we build a dataset for OV-MER, we also offer a complete description of the data construction process in Section 2 and Appendix. In summary, we have made every effort to ensure its reproducibility.
Appendix A Related Work

### v2

@@ 3 Evaluation Metric
3.1 Grouping
We propose two grouping strategies: one based on GPT and the other based on the emotion wheel (EW) [20]. In the experiments, we use GPT-based grouping by default.
GPT-based Grouping.
The most direct approach is to use GPT-3.5 to group all labels based on their similarity: Please assume the role of an expert in the field of emotions. We provide a set of emotions. Please group the emotions, with each group containing emotions with the same meaning. Directly output the results. The output format should be a list containing multiple lists. However, the evaluation results may be affected by the API version. For example, if OpenAI deprecates an old API, the results based on that API will become difficult to reproduce. Additionally, this process is costly (see Appendix N). Therefore, we attempt to find a replacement for GPT-based grouping.
EW-based Grouping.
EW is a psychological model that categorizes emotions in a structured manner. The inner part shows core emotions while moving to the outer part reveals more nuanced emotions. Therefore, EW naturally provides emotion grouping information. Since there is no consensus on EW, we select five typical wheels (see Appendix O).
Before calculating the metrics, we define some symbols. We group the labels by their levels from the innermost to the outermost as $L_{w_{1}}^{1}$, $L_{w_{1}}^{2}$, and $L_{w_{1}}^{3}$. Next, we define a function $m_{w_{1}}^{i\rightarrow j}(\cdot)$ that maps the labels in $L_{w_{1}}^{i}$ to the corresponding labels in $L_{w_{1}}^{j}$. From inner to outer ($i<j$), $m_{w_{1}}^{i\rightarrow j}(\cdot)$ is a many-to-one mapping; from outer to inner ($i>j$), $m_{w_{1}}^{i\rightarrow j}(\cdot)$ is a one-to-many mapping. We collect all the labels from these emotion wheels and represent them as EW, i.e., $\{L_{w_{i}}^{j},1\leq i\leq 5,1\leq j\leq 3\}$. We denote the labels in EW as $y_{w}$.
[…]
(1) M1. We use $g^{\prime}(\cdot)$ to map each label to its $y_{f}$.
(2) M2. We use $f^{\prime}(g^{\prime}(\cdot))$ to map each label to its $y_{w}$.
(3) M3. We use the emotion wheel during metric calculation. Specifically, we first use $f^{\prime}(g^{\prime}(\cdot))$ to map each label to its $y_{w}$. Then, we define two grouping functions, L1 and L2. For L1, we map all labels to their corresponding $L_{w_{i}}^{1}$:
For L2, we map all labels to their corresponding $L_{w_{i}}^{2}$:
3.2 Metric Definition
Then, we convert the above emotion grouping information into a function $G(\cdot)$, which can map each label to its group ID. Specifically, suppose $\{y_{i}\}_{i=1}^{M}$ and $\{\hat{y}_{i}\}_{i=1}^{N}$ are the ground truth and predictions, where $M$ and $N$ are the number of labels. We first map each label into its group ID: $\mathcal{Y}=\{G(x)|x\in\{y_{i}\}_{i=1}^{M}\}$ and $\hat{\mathcal{Y}}=\{G(x)|x\in\{\hat{y}_{i}\}_{i=1}^{N}\}$. Then, we design set-based metrics for performance evaluation. Specifically, $\mbox{Precision}_{\mbox{s}}$ indicates the number of correctly predicted labels; $\mbox{Recall}_{\mbox{s}}$ indicates whether the prediction covers all ground truth:
$\mbox{Precision}_{\mbox{s}}=\frac{|\mathcal{Y}\cap\hat{\mathcal{Y}}|}{|\hat{\mathcal{Y}}|},\;\mbox{Recall}_{\mbox{s}}=\frac{|\mathcal{Y}\cap\hat{\mathcal{Y}}|}{|\mathcal{Y}|}.$ (3)
$\mbox{F}_{\mbox{s}}$ is the harmonic mean of two metrics, which is used for the final ranking:
$\mbox{F}_{\mbox{s}}=2\times\frac{\mbox{Precision}_{\mbox{s}}\times\mbox{Recall}_{\mbox{s}}}{\mbox{Precision}_{\mbox{s}}+\mbox{Recall}_{\mbox{s}}}.$ (4)
It is important to note that changing the label order in $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ does not result in any score change.
4 Baselines for OV-MER
4.1 CLUE Generation
Figure 2 illustrates the generation process of CLUE-Multi, where we combine text with checked visual and acoustic clues. In this section, we further introduce some variants.
CLUE-A/T/V.
To reveal the modality impact, we propose three variants of CLUE-Multi: CLUE-Audio, CLUE-Text, and CLUE-Video. In Figure 3, we illustrate their generation process. (1) CLUE-Audio: We observe that ALLM cannot fully leverage the text, and using an additional LLM to emphasize the text can further improve performance, which is also verified in Section 5. Therefore, we merge the checked acoustic clues with text using an additional LLM; (2) CLUE-Text: We only use the text to infer emotional states; (3) CLUE-Video: Since the visual content does not contain audio and text, we only use the checked visual clues. See Appendix P for more examples.
CLUE-MLLM.
MLLMs can address various multimodal tasks. Since emotion recognition relies on temporal information, we choose models that support at least video or audio. To generate CLUE-MLLM, we first use ALLM or VLLM to extract emotion-related descriptions, and then combine these descriptions with text using LLM. Compared with CLUE-Multi, this process does not use manually checked clues. Appendix Q provides model cards and relevant prompts. This paper aims to build a zero-shot benchmark for OV-MER, without the training process. All models are implemented in PyTorch, and all inference processes are carried out using a 32G NVIDIA Tesla V100 GPU.
@@ 4.1 CLUE Generation
4.2 Metric Calculation
5 Results and Discussion
In this section, we default to using GPT-based grouping and employ GPT-3.5 (“gpt-3.5-turbo-16k-0613”) as LLM. We generally report evaluation results in both languages, but if no specific language is mentioned, we default to reporting results for the English branch. To mitigate the impact of randomness, we conduct each experiment twice and report the average scores and standard deviations.
Main Results on CLUE-M/A/T/V.
For CLUE-M/A/T/V, most baselines use manually checked clues, serving as performance upper bounds of different modality combinations. In Table 2, we observe that CLUE-Multi performs the best, highlighting the importance of multimodal information in MER. In contrast, CLUE-Text performs the worst. This is because our OV-MERD dataset is derived from MER2023, where the contribution of text is smaller compared to audio and video [21]. Relying solely on text makes it difficult to recognize emotions accurately.
Main Results on CLUE-MLLM.
In Table 2, we introduce a heuristic baseline called Random, where we randomly select a label from basic emotions. This baseline reflects the lower bound of performance. We observe that MLLM generally outperforms Random, indicating that MLLM can partially address OV-MER. However, the performance of MLLM remains unsatisfactory, highlighting the limitations of existing MLLMs and the challenges of OV-MER. Furthermore, models that perform well in Chinese often perform well in English, suggesting that the impact of language differences on rankings is limited.
(a) Length (H)
(b) #Label (H)
(c) Word cloud (H)
(d) Length (H+L)
(e) #Label (H+L)
(f) Word cloud (H+L)
Figure 5: Human-only (H) vs. Human-LLM (H+L) strategy.
Model GPT M1 M2 M3-W1 M3-W2 M3-W3 M3-W4 M3-W5
Qwen-Audio 38.13±0.05 20.49±0.01 23.37±0.01 43.85±0.03 41.52±0.28 39.46±0.28 36.64±0.03 35.89±0.08
Otter 43.51±0.09 22.21±0.06 27.99±0.02 49.75±0.11 49.93±0.11 51.03±0.04 47.54±0.00 50.51±0.03
Video-LLaMA 44.73±0.14 23.56±0.08 28.39±0.17 52.90±0.12 53.60±0.04 47.50±0.20 52.97±0.09 46.39±0.12
VideoChat 45.53±0.11 22.15±0.00 26.24±0.08 47.79±0.07 47.76±0.11 46.78±0.11 49.53±0.15 45.93±0.18
SECap 45.72±0.09 26.52±0.01 32.88±0.03 52.26±0.03 52.11±0.03 50.77±0.03 50.43±0.03 49.97±0.03
Video-LLaVA 47.07±0.16 25.47±0.12 30.73±0.11 54.65±0.10 54.54±0.02 52.29±0.05 52.45±0.06 52.97±0.10
SALMONN 47.96±0.04 23.57±0.02 28.83±0.03 54.90±0.15 54.29±0.06 56.25±0.01 50.53±0.09 53.65±0.04
VideoChat2 49.07±0.26 26.92±0.09 31.40±0.10 52.38±0.13 53.56±0.13 52.14±0.23 50.63±0.19 51.37±0.14
Video-ChatGPT 50.52±0.06 28.99±0.04 34.05±0.05 57.66±0.04 57.37±0.00 55.50±0.13 55.24±0.02 52.93±0.05
LLaMA-VID 51.25±0.09 28.28±0.04 32.85±0.03 56.59±0.04 57.49±0.03 55.12±0.05 56.62±0.15 53.03±0.08
mPLUG-Owl 52.73±0.13 27.47±0.17 32.47±0.19 57.60±0.23 56.32±0.26 55.67±0.19 55.06±0.17 54.44±0.13
Chat-UniVi 53.08±0.01 28.89±0.02 33.23±0.08 57.00±0.06 57.50±0.03 56.80±0.01 55.86±0.07 55.81±0.02
PCC score 0.887 0.857 0.911 0.913 0.904 0.899 0.885
@@ 5 Results and Discussion
To verify the effectiveness of our human-LLM strategy, we additionally introduce a baseline using human-only annotation. In Figure 5, we compare two strategies from three aspects: the length distribution of generated descriptions, the distribution of sample-wise label numbers, and the word cloud. We observe that through human-LLM collaboration, we can obtain longer descriptions, provide more diverse labels for each sample, and generate a broader range of emotions. These results show that human-only annotation generally focuses on primary emotions while neglecting minor ones. With the pre-annotation and semantic reasoning capabilities of LLMs, we can obtain richer emotional labels. These results validate the effectiveness of our human-LLM collaborative strategy. Meanwhile, these results suggest that the LLM-driven approach does not lead to a narrow or biased interpretation of emotions, but rather helps uncover more subtle emotional nuances. We provide additional analysis in the Appendix R.
Figure 6: Ablation.
@@
We reveal the impact of different CLUE-MLLM generation strategies. Figure 6 introduces three methods: 1) S0 does not use text and inputs the video into MLLM; 2) S1 inputs both text and video into MLLM; 3) S2 first uses MLLM to extract descriptions and then combines with text using another LLM, same with the strategy in Figure 3. In Figure 4, S1 and S2 generally outperform S0, indicating the importance of the text content in OV-MER. Moreover, S2 typically performs better than S1. The reason is that inputting video and text into the MLLM simultaneously increases the task difficulty, and current MLLMs may struggle to handle complex prompts. S2 divides this process into two steps, reducing task complexity and achieving better performance. Therefore, we adopt S2 as the default strategy. More results are provided in Appendix Q.
@@
We propose two grouping strategies: GPT-based and EW-based grouping. In this section, we explore the relationship between them. Table 3 reports $\mbox{F}_{\mbox{s}}$ for different EW-based strategies. For M3, we report the scores of the L1 strategy by default, and the complete results are provided in Appendix S. We observe that the PCC scores between GPT-based and EW-based groupings are relatively high, indicating that EW-based metrics can serve as an alternative to GPT-based metrics.
@@
GPT-based vs. Matching-based Metrics.
In Table 2, CLUE-Multi performs the best. This leads to a hypothesis: Do sentences that are more “similar” to CLUE-Multi yield better emotion recognition performance? The most common way to measure “similarity” is through matching-based metrics, with BLEU1, BLEU4, METEOR, and ROUGE1 being the most widely used. Therefore, we use CLUE-MLLM as input and calculate both GPT-based and matching-based metrics (see Figure 7(a)), followed by calculating their PCC scores (see Figure 7(b)). From the experimental results, we have some interesting observations. First, the same metric across different languages typically exhibits high correlations. However, the correlation between GPT-based and matching-based metrics is not strong. For example, the highest PCC score between “F(E)” and matching-based metrics is 0.77. These results highlight the limitations of using matching-based metrics to evaluate the OV-MER task. The main reason is that matching-based metrics focus on low-level word-level matches, while emotion understanding is a relatively complex and high-level perceptual task. In Appendix T, we provide more examples for clarification.
Informative Comparison.
In this section, we conduct a user study to verify whether our dataset provides greater informativeness than basic emotions. Specifically, we hired four annotators and randomly selected samples from our dataset. For each sample, we provided both the basic emotion label and the OV-MERD label. The following instruction was given: Which label provides greater informativeness? The label with more information was marked as 1, and the other label as 0. Experimental results demonstrate that 97.5% of the results believed that our OV-MERD provided more information than basic emotions, verifying the effectiveness of our OV-MERD in emotion representation.
@@ 6 Conclusion
Appendix A Task Comparison
There are various MER tasks. In this section, we provide an in-depth analysis of the differences between our proposed OV-MER and existing tasks.
One-Hot MER (OH-MER) is most widely discussed in the community. In this task, each sample is labeled with only one emotion. We should design a framework to predict the most likely label from a predefined emotional taxonomy [22]. Current research primarily focuses on the framework design and multimodal fusion strategy. For the former, research has shifted from traditional classifiers (e.g., SVM [23]) to deep models (e.g., Transformers [24]). For the latter, researchers mainly focus on how to align heterogeneous features (e.g., word alignment [25] and implicit alignment [26]) for subsequent multimodal fusion.
Multi-Label MER (ML-MER) considers the complexity of emotions (i.e., multiple emotions often occur simultaneously) and allows the model to predict multiple labels. The most straightforward solution is to use binary classifiers for each emotion [27, 28]. However, this approach is based on the assumption of emotion independence, ignoring the correlations between different emotions. Therefore, [29] and [30] proposed a sequence-to-emotion model, transforming the multi-label emotion recognition task into an emotion sequence prediction task, to further consider the correlations between emotions.
Partial Label Learning MER (PLL-MER) differs from the tasks mentioned above. In PLL-MER, each sample is associated with a set of candidate labels, only one of which is correct [31]. Research on PLL-MER can be roughly divided into two categories: average-based methods and identification-based approaches. The former assumes that each candidate label has an equal probability of being the ground truth [32], while the latter directly identifies the ground truth and maximizes its estimated probability [33].
Fine-Grained MER (FG-MER) differs from the above tasks in the label space. Previous tasks typically rely on coarse-grained emotion taxonomies, such as the widely accepted basic emotion taxonomies, like Ekman [6] or Plutchik [20] emotions. Differently, FG-MER aims to use fine-grained emotion taxonomies to capture more subtle emotions [34]. After expanding the label space, FG-MER follows the typical solution of OH-MER.
OV-MER shares some similarities with existing tasks, such as allowing the prediction of multiple labels, like ML-MER. However, OV-MER is fundamentally different from previous tasks:
  • Task Definition. The main distinction of OV-MER from other tasks lies in its focus on generalizing to unseen or new labels, whereas other tasks operate within a predefined taxonomy, either coarse-grained or fine-grained taxonomies. This is also the reason why we use the term “open” in the task name.
  • Emotional Complexity. Human emotional states are diverse and nuanced. As psychologist Plutchik pointed out, humans can express approximately 34,000 different emotions [20]. Predefined taxonomies that categorize the full spectrum of emotions into a limited set of labels inevitably overlook some subtle emotional states. In contrast, OV-MER allows the model to understand and predict any emotion, enabling a more accurate modeling of the complex nature of human emotions.
  • Relationship Between OV-MER and ML-MER. Theoretically, ML-MER can be converted to OV-MER by spanning the label space to a complete set that includes all the emotion labels in the language (e.g., around 34,000 different emotions [20]). However, it is not feasible in practice to construct such a kind of dataset, i.e., annotators need to label each sample with 34,000 different emotions, and it’s hard to collect sufficient samples for every emotion category, let alone multiple labels).
  • Evaluation Metrics. In OV-MER, we can use any emotion word to describe a person’s emotional state. Therefore, the test set may contain new emotion labels that are not seen during training. In contrast, traditional methods require the label space in both training and test sets to be strictly consistent. This is why we propose new evaluation metrics for OV-MER.
  • Solution. Previous tasks rely on predefined label spaces, meaning that the predicted output is a fixed-size $M$-dimensional vector, where $M$ is the size of the label space. Therefore, previous tasks can be solved using discriminative methods. However, in OV-MER, we do not constrain the label space, making discriminative methods unsuitable. Therefore, we adopt a generative approach, leveraging the rich output vocabulary of LLMs to construct our solution.
Appendix B Limitations
Firstly, the main contribution of this paper is the definition of a new task and the conduct of foundational research. In the future, we plan to design more effective frameworks to solve OV-MER. Specifically, we will incorporate more emotion-related instruction datasets to finetune MLLMs, thereby enhancing their emotion recognition ability. Meanwhile, how to integrate subtitle information and fuse multimodal inputs also plays a crucial role. We will also consider these aspects in the framework design.
Secondly, we evaluate some MLLMs, but not all models are covered. In the future, we will expand the scope of evaluation to cover more MLLMs to enrich our benchmark.
Thirdly, this paper does not involve cultural differences. Specifically, our original data is in Chinese, and the annotators we hired are also native Chinese speakers. In the future, we will also try to extend our method to other cultures and further analyze cultural differences.
Appendix C Detailed Motivation
Video Emotion vs. Facial Emotion.
Video emotion is more complex than facial emotion. This is because, in videos, we need to capture subtle changes in the temporal dimension and integrate multimodal clues. Take Figure 1(b) as an example. In the temporal dimension, we need to infer a person’s nervousness based on his stuttering; in the multimodal dimension, we need to combine information from different modalities to gain a more comprehensive understanding of emotion. Due to the complexity of video emotion, using a single label is limiting, and more discrete labels are required to better describe video emotion. This is also the motivation behind our OV-MER task.
Label Importance.
In OV-MER, we do not assign different levels of importance to each label. Every emotion holds equal significance, and neglecting anyone can impact the performance of downstream tasks. For example, if a human-computer interaction system overlooks any emotion, it may fail to generate appropriate responses.
Appendix D Related Work